# Request a Note: How the Request Function Shapes X’s Community Notes System

### v1

@@ 3.2.1. Sentiments
Sentiments are an important factor that influences the dissemination of online content (Chuai and Zhao, 2022; Robertson et al., 2023). We extract sentiments in the posts content using a state-of-the-art sentiment classification model published on HuggingFace. The predicted sentiment categories include positive, neutral, and negative. Specifically, we compute the probabilities of positive and negative sentiments for each collected post using the TwitterroBERTa-base model (2022 updated). It was trained on 124 million posts created from January 2018 to December 2021, and fine-tuned for sentiment analysis. This model achieves superior predictive performance compared to other similar sentiment models (Loureiro et al., 2022). We denote the probabilities of positive sentiment and negative sentiment in each post as $\mathit{Positive}$ and $\mathit{Negative}$, respectively. On average, requested posts exhibit significantly higher negative sentiment (mean of $0.460$) than positive sentiment (mean of $0.131$; $t=$ $199.208$, $p<$ $0.001$).
@@ 3.2.3. Content types and misleadingness
We distinguish between factual claims, which can be independently verified, and personal opinions, which are more subjective content to express users’ personal feelings. This distinction is important for understanding how requests and community notes may prioritize verifiable content over opinionated posts. To operationalize this, we use GPT-4.1 to classify requested posts as factual claims or personal opinions. In addition, prior work suggests that LLMs can achieve a high performance in identifying misleading information, which is comparable to human annotators (Choi and Ferrara, 2024; DeVerna et al., 2024; Costabile et al., 2025). Building on this, we employ GPT-4.1 to assess the misleadingness of collected posts. We provide task-specific prompts to the model and manually validate its outputs through a subset of random posts to ensure reliability (see details in Suppl. S1). This approach allows us to analyze both content types and the likelihood of misinformation in the requested posts. The requested posts are more likely to be claims (mean of 0.666), compared to opinions (mean of 0.522; $t=$ $84.797$, $p<$ $0.001$). Additionally, the GPT-estimated misleadingness is, on average, $0.428$ in the requested posts.
@@ 4.1. Likelihood of Receiving Community Notes
Of all the requested posts, $52,862$ (53.6%) received at least one community note. To investigate the factors associated with this outcome (RQ1), we estimate a logistic regression model predicting the likelihood that a requested post receives a community note. The estimation results are presented in Fig. 2.
@@
Post content characteristics. We analyze how post content characteristics shape the likelihood of receiving community notes. For sentiments, we find that posts expressing stronger negative sentiment are significantly less likely to receive community notes (coef.=0.077coef.=-0.077, $p<0.01$). For topics, we find that posts related to Politics (coef.=0.334coef.=-0.334, $p<0.001$) and Science & Technology (coef.=0.117coef.=-0.117, $p<0.001$) are less likely to receive community notes compared to posts not associated with these topics. For example, political posts have 28.4% lower odds of receiving community notes compared to non-political posts. In contrast, posts related to Entertainment ($coef.=0.086$, $p<0.001$) and Finance & Business ($coef.=0.073$, $p<0.001$) are more likely to receive community notes compared to posts not associated with these topics. Additionally, posts containing media are more likely to receive community notes compared to posts without media ($coef.=0.205$, $p<0.001$), corresponding to a 22.7% increase in the odds of receiving community notes. For claims vs. opinions, we find that the probability that a post contains a claim is not significantly associated with its likelihood of receiving a community note ($coef.=0.037$, $p=0.271$), whereas posts expressing opinions are significantly less likely to receive notes (coef.=0.096coef.=-0.096, $p<0.001$). Importantly, we find that posts with higher GPT-estimated misleadingness are significantly more likely to receive community notes ($coef.=0.308$, $p<0.001$). This suggests that Community Notes contributors tend to prioritize posts with a higher potential for spreading misinformation.
@@
Post author characteristics. We further evaluate how post author characteristics shape the likelihood of receiving community notes. For verified types of post authors, we find that posts from blue (coef.=0.122coef.=-0.122, $p<0.001$) and business (coef.=0.203coef.=-0.203, $p<0.001$) accounts are less likely to receive community notes than posts from unverified accounts. Additionally, posts from accounts with a higher number of followees are less likely to receive community notes (coef.=0.060coef.=-0.060, $p<0.001$). In contrast, we find that posts from accounts with higher misinformation exposure scores ($coef.=0.621$, $p<0.001$) or right-leaning partisanship ($coef.=0.053$, $p<0.01$) are more likely to receive community notes. This finding suggests that contributors prioritize posts from accounts with higher risk of spreading misinformation and right-leaning partisanship.
@@
In summary, the likelihood that a requested post receives a community note is shaped by both posts’ and their authors’ characteristics. Requested posts with negative sentiment, opinionated framing, or authored by blue/business accounts are less likely to attract notes, as are those related to politics or science. In contrast, posts related to entertainment or finance, those containing media, and posts flagged by GPT as more misleading are significantly more likely to receive notes. At the author level, posts from accounts with higher misinformation exposure or right-leaning partisanship are also more likely to be annotated. These findings suggest that Community Notes contributors who write notes and requestors who submit requests may have different selection patterns that are associated with both post features and author attributes.
@@ 4.2. Request Timing and Request-Fostered Notes
Request timing relative to note creation. As shown in Fig. 3(a), the median time from post creation to the submission of first request is 2.72 hours, while it takes significantly longer for a post to accumulate five requests (median of 13.4 hours; $p_{MWU}<0.001$). We find that the response time of community notes (median of 8.2 hours) is significantly shorter than the time it typically takes for posts to receive five requests (Fig. 3(b); $p_{MWU}<0.001$). Furthermore, Fig. 3(c) shows that top writers (median of 7.78 hours) generate notes almost an hour faster than other contributors (median of 8.76 hours; $p_{MWU}<0.001$). Taken together, these findings suggest that community notes tend to be written before the fifth request is submitted (i. e., request threshold)—implying that requests may not be the primary driver of note creation. Consistent with the guidelines of the request feature, we find that only an estimated 12.1% of requested posts were likely influenced by the request activity, as their associated notes were written by top writers after the request threshold was reached (Fig. 3(d)). Therefore, the effect of the request feature in driving note creation is still limited.
@@
Request-fostered notes from top writers. We further employ logistic regression to compare how top writers select posts after seeing request alerts versus when writing notes without such alerts. The estimation results are reported in Fig. 4. The coefficient estimates for Politics ($coef.=0.247$, $p<0.001$) and Science & Technology ($coef.=0.091$, $p<0.05$) are significantly positive. This suggests that top writers select more political and scientific posts when receiving request alerts. On the contrary, the coefficient estimates for Entertainment (coef.=0.097coef.=-0.097, $p<0.01$) and opinionated content (coef.=0.102coef.=-0.102, $p<0.05$) are significantly negative. This suggests that top writers choose less entertainment-related and opinionated content after receiving request alerts. These results suggest that requests can shift contributors’ attention toward political and scientific content, while diverting attention away from opinionated posts or those related to entertainment. In particular, the coefficient estimate for GPT-estimated misleadingness is significantly positive ($coef.=0.451$, $p<0.001$), suggesting that top writers prioritize posts with higher potential misleadingness when guided by request alerts.
In summary, out of all requested posts eligible for community evaluation, only 12.1% appear to have been generated by top writers through request alerts. Furthermore, requests typically arrive later than community notes themselves, as the time from post creation to request alerts is significantly longer than the time from post creation to note generation, limiting the request feature’s potential to foster note writing. However, by comparing notes written by top writers before and after the request threshold, we find that requests shift contributors’ attention toward political content while prioritizing posts with higher potential misleadingness. Thus, although the overall effect of requests on note creation remains modest, the feature shows promise in channeling contributor attention toward high-stakes content at greater risk of misinformation.
@@ 4.3. Algorithmic Evaluation of Community Notes
To address RQ3, we downloaded and rerun the source code of the note selection algorithm released by $\mathbb{X}$. Fig. 5(a) presents the distribution of note intercepts and note factors from the note selection algorithm. The note intercepts reflect the perceived helpfulness of community notes across raters, with higher intercepts indicating higher helpfulness. The note factors capture the polarization in ratings, with values deviating from 0 indicating greater disagreement among raters regarding the helpfulness of community notes. Here, we categorize all community notes into three groups: writer-only (notes on posts that received fewer than five requests and were solely written by contributors), request-fostered (notes authored by top writers after the fifth request, likely influenced by requests), and request-related (all remaining notes on posts that received five or more requests but not classified as “request-fostered”). We then analyze the helpfulness and polarization (measured as the absolute value of note factors) of community notes across these categories.
@@
Figure 5. Overview of note evaluations and source domains in community notes. (a) The distribution of estimated note intercepts (helpfulness) and estimated note factors (polarization) from the note selection algorithm. (b) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and writer-only notes. (c) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-fostered notes and writer-only notes. (d) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and request-fostered notes. (d) The violin plots showing the distributions of domain bias in community notes across the three categories: writer-only, request-related, and request-fostered. (e) The violin plots showing the distributions of domain quality in community notes across the three categories: writer-only, request-related, and request-fostered.
[…]
@@
Overview of note evaluations and source domains in community notes. (a) The distribution of estimated note intercepts (helpfulness) and estimated note factors (polarization) from the note selection algorithm. (b) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and writer-only notes. (c) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-fostered notes and writer-only notes. (d) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and request-fostered notes. (d) The violin plots showing the distributions of domain bias in community notes across the three categories: writer-only, request-related, and request-fostered. (e) The violin plots showing the distributions of domain quality in community notes across the three categories: writer-only, request-related, and request-fostered.
@@
Taken together, these findings indicate that, although requests in general are associated with less helpful and more polarized notes, request-fostered notes written by top writers tend to achieve higher quality—being both more helpful and less polarized than request-related notes. To further substantiate this pattern, we estimate a logistic regression predicting whether a note is request-fostered versus request-related using helpfulness and polarization as predictors. The results show a strong positive association for helpfulness ($coef.=2.528$, $p<0.001$) and a significant negative association for polarization (coef.=0.231coef.=-0.231, $p<0.001$), confirming that request-fostered notes stand out within the request context by being both more helpful and less polarized than other notes. Subsequently, we investigate what factors could explain this quality advantage.
@@
Because Community Notes contributors are anonymized, rater bias toward specific writers is unlikely. Observed differences in helpfulness or polarization likely reflect the types of posts and authors contributors choose to fact-check. We therefore examine which notes, posts, and authors are associated with higher helpfulness or polarization.
Factors associated with note helpfulness. The estimation results from a linear regression model predicting note helpfulness are shown in Fig. 6. We find that the unbiasedness (i. e., neither left-leaning nor right-leaning) of external domains cited in community notes is positively associated with note helpfulness ($coef.=0.005$, $p<0.05$), suggesting that notes referencing politically neutral sources are perceived as more helpful compared to those citing left- or right-leaning domains. In contrast, the information quality of external domains is not significantly associated with helpfulness ($coef.=0.015$, $p=0.080$).
Regarding post content, notes on posts with higher negative sentiment are rated as less helpful (coef.=0.013coef.=-0.013, $p<0.01$). Posts related to Politics (coef.=0.015coef.=-0.015, $p<0.001$) and Science & Technology (coef.=0.012coef.=-0.012, $p<0.001$) tend to receive less-helpful notes compared to posts without these topics. Additionally, posts with higher opinion scores are associated with lower note helpfulness (coef.=0.023coef.=-0.023, $p<0.001$). Conversely, community notes on posts containing media ($coef.=0.009$, $p<0.001$) or with higher estimated misleadingness ($coef.=0.014$, $p<0.01$) are rated as more helpful. With respect to post authors, notes on posts from the blue (coef.=0.010coef.=-0.010, $p<0.01$) and government (coef.=0.013coef.=-0.013, $p<0.01$) accounts, as well as from authors with more followers (coef.=0.027coef.=-0.027, $p<0.001$), are less helpful compared to those from unverified accounts with fewer followers. In contrast, higher misinformation exposure scores of post authors are positively associated with note helpfulness ($coef.=0.091$, $p<0.001$).
In summary, the helpfulness of community notes is positively associated with (i) citations to unbiased domains, (ii) posts containing media, (iii) post misleadingness, and (iv) misinformation exposure scores of post authors. On the contrary, the helpfulness of community notes is negatively associated with (i) negative sentiment in posts, (ii) political and scientific topics, (iii) opinionated content, (iv) blue and government accounts, and (v) the number of followers.

### v2

@@ 3.2.1. Sentiments
Sentiments are an important factor that influences the dissemination of online content (Chuai and Zhao, 2022; Robertson et al., 2023). We extract sentiments in the posts content using a state-of-the-art Twitter-roBERTa-base sentiment classification model (2022 updated), which is available on HuggingFace. [^2] https://huggingface.co/cardiffnlp/twitter-roberta-base-sentiment-latest This model was trained on 124 million posts created from January 2018 to December 2021, and fine-tuned for sentiment analysis. It achieves superior predictive performance on the TweetEval benchmark, with a reported macro-recall of 0.737 on the sentiment classification task (Loureiro et al., 2022). The model outputs probability estimates for three sentiment categories: positive, neutral, and negative, with the sum of the probabilities across these categories equaling to 1. On average, requested posts exhibit strongest negative sentiment (mean of $0.460$), followed by neutral sentiment (mean of $0.409$; $t(\mathit{negative},\mathit{neutral})=$ $28.651$, $p<$ $0.001$) and positive sentiment (mean of $0.131$; $t(\mathit{negative},\mathit{positive})=$ $199.208$, $p<$ $0.001$). In this study, we focus on affectively charged content; thus, to avoid multicollinearity in our empirical regression analyses, we retain the probabilities of positive and negative sentiments for each collected post and omit the neutral probability.
@@ 3.2.3. Content types and misleadingness
Traditional expert-based fact-checking and emerging LLM-based misinformation detection methods primarily aim to identify checkworthy, objective, and verifiable claims, subsequently evaluating their truthfulness (Chuai et al., 2025b; Costabile et al., 2025). These approaches focus on factual claims and often overlook opinion-based content that may contain misleading interpretations (Ni et al., 2024). In contrast, the Community Notes system relies on crowd consensus to assess the misleadingness of selected posts and surface community fact-checks, with the potential to address both factual inaccuracies and misleading opinions (see the misleading criteria in Suppl. S1). Given this, we distinguish between factual claims, which explicitly presents some verifiable facts, and personal opinions, which are more subjective content to express users’ personal judgments (see full definitions in Supp. S2). This distinction is important for understanding how requests and community notes select verifiable content and opinionated posts. Notably, recent critiques also highlight that a substantial portion of community notes have been attached to opinion-based posts and published for debate rather than moderation (Razuvayevskaya et al., 2025). To assess whether requestors and contributors target content that is plausibly misleading, rather than merely content they disagree with, we analyze the misleadingness of requested posts.
Given that recent work demonstrates the comparable performance of LLMs, such as GPT, to human evaluations in text quality and their potential applications in large-scale empirical analysis (Chiang and Lee, 2023; Rathje et al., 2024; DeVerna et al., 2024), we use GPT-4.1, a state-of-the-art LLM model at the time of this study, to classify content types (claims vs. opinions) and to estimate the misleadingness of requested posts. To ensure that the model’s outputs align with the scope of our analysis, we provide GPT a task-specific prompt incorporating the definitions and criteria for misleading information, factual claims, and opinions (see full prompt in Suppl. S2). The GPT-estimated scores for claim, opinion, and misleadingness for each post range from 0 to 1, with higher values indicating a stronger presence of the corresponding attribute. Additionally, a post can contain both claims and opinions, and their estimated scores are not required to sum to 1.
To assess the reliability of GPT-generated annotations, we first randomly select 200 posts from the requested dataset and conduct a manual validation using a small crowd. Specifically, seventeen university-level student assistants are recruited and trained on the relevant definitions and criteria to annotate whether each assigned posts contain claims, opinions, and misleading information. Each post is independently reviewed by a minimum of six assistants to reduce individual-level noise and ensure robust aggregation, with final annotations determined by averaging their ratings. For comparison, we transform the GPT-estimated scores into binary classes using a threshold of 0.5, and compare the GPT-generated labels with the assistant-annotated labels. We find that the GPT-based approach demonstrates strong performances for claims ($\mathit{Accuracy}=$ 0.840, $\mathit{Weighted~F1}=$ 0.838), opinions ($\mathit{Accuracy}=$ 0.805, $\mathit{Weighted~F1}=$ 0.810), and misleadingness ($\mathit{Accuracy}=$ 0.800, $\mathit{Weighted~F1}=$ 0.800). These results suggest that the GPT-generated outputs are comparable to the aggregated human judgments, supporting their use in our large-scale analysis (see its validation with expert fact-checks from PolitiFact in Suppl. S3). Overall, the requested posts are more likely to contain claims (mean of 0.687), compared to opinions (mean of 0.489; $t=$ $111.269$, $p<$ $0.001$). Additionally, the GPT-estimated misleadingness is, on average, $0.398$ in the requested posts. [^3] We use GPT to annotate the 200 sampled posts twice and assess the consistency of its outputs across the two rounds. The outputs show strong agreement across all three variables: claims (Cohen’s $\kappa=0.920$), opinions (Cohen’s $\kappa=0.828$), and misleadingness (Cohen’s $\kappa=0.816$). In addition, we further validate the reliability of GPT-based approach using expert fact-checks from PolitiFact ($\mathit{Accuracy}=$ 0.801, $\mathit{Weighted~F1}=$ 0.857; see details in Suppl. S3).
@@ 3.6. Replication of Note Selection Algorithm
3.7. Estimation Models
In this study, we employ a series of regression models and conduct empirical analyses using our observational dataset and pre-defined variables to address the research questions.
For RQ1 and RQ2, we specify a logistic regression model to examine how contributors select posts for fact-checking:
(1) $logit(P(\mathit{Noted_{i}}))=$ $\beta_{0}+\bm{\beta_{1}^{\prime}\mathit{X_{i}^{Post}}}+\bm{\beta_{2}^{\prime}\mathit{X_{i}^{Poster}}},$
where $P(\mathit{Noted_{i}})$ indicates the probability of the post $i$ receiving a community note, and $\beta_{0}$ is the intercept. The vector $\bm{\mathit{X_{i}^{Post}}}$ contains variables describing characteristics of post content, including sentiments, topics, content types, misleadingness, and media. Additionally, the word count of post $i$ is included to control for the post length (Chuai et al., 2024b; Robertson et al., 2023). Topics and media are represented as dummy variables, while all other post-level features are treated as continuous variables. The vector $\bm{\mathit{X_{i}^{Poster}}}$ contains variables describing characteristics of post authors, including account types, account age, followers, followees, misinformation exposure scores, and partisan scores. Account types are encoded as dummy variables, while all other author-level features are treated as continuous variables. The vectors $\bm{\beta_{1}}$ and $\bm{\beta_{2}}$ represent the coefficients of the post-level and author-level variables, respectively, and quantify their effects on the likelihood of receiving community notes. Since logistic regression coefficients are expressed in log-odds, we exponentiate them into odds ratios to facilitate interpretation and address our RQ1 and RQ2.
For RQ3, we specify a linear regression model to examine how the extracted features of community notes, posts, and post authors relate to the helpfulness and polarization of community notes:
(2) $Y_{j}=$ $\beta_{0}+\bm{\beta_{1}^{\prime}\mathit{X_{j}^{Note}}}+\bm{\beta_{2}^{\prime}\mathit{X_{j}^{Post}}}+\bm{\beta_{3}^{\prime}\mathit{X_{j}^{Poster}}}+\epsilon_{j},$
where $Y_{j}$ represents either the helpfulness score or the polarization score of note $j$. $\beta_{0}$ is the intercept, and $\epsilon_{j}$ is the residual term for the linear regression. The vector $\bm{\mathit{X_{j}^{Note}}}$ includes note-specific characteristics, including domain quality and domain unbiasedness. They are treated as continuous variables. The vectors $\bm{\mathit{X_{j}^{Post}}}$ and $\bm{\mathit{X_{j}^{Poster}}}$ contains post-level and author-level features, respectively, and remain the same as defined in Eq. 1. Coefficient estimates for $\bm{\beta_{1}}$, $\bm{\beta_{2}}$, and $\bm{\beta_{3}}$ reflect how note-, post-, and author-level attributes are associated with variations in note helpfulness and polarization, thus allowing us to assess RQ3.
@@ 4.1. Likelihood of Receiving Community Notes
Of all the requested posts, $52,862$ (53.6%) received at least one community note at the time of data collection. To investigate the factors associated with this outcome (RQ1), we estimate a logistic regression model predicting the likelihood that a requested post receives a community note. The estimation results are presented in Fig. 2.
@@
Post content characteristics. We analyze how post content characteristics shape the likelihood of receiving community notes. For sentiments, we find that posts expressing stronger negative sentiment are significantly less likely to receive community notes ($\mathit{coef.}=-0.095$, $\mathit{odds~ratio}=0.909$, $p<0.01$). For instance, posts classified as purely negative exhibit 9.1% lower odds of receiving notes compared with non-negative posts. For topics, we find that posts related to Politics ($\mathit{coef.}=-0.337$, $\mathit{odds~ratio}=0.714$, $p<0.001$) and Science & Technology ($\mathit{coef.}=-0.116$, $\mathit{odds~ratio}=0.890$, $p<0.001$) are less likely to receive community notes compared to posts not associated with these topics. For example, political posts have 28.6% lower odds of receiving community notes compared to non-political posts. In contrast, posts related to Entertainment ($\mathit{coef.}=0.084$, $\mathit{odds~ratio}=1.088$, $p<0.001$) and Finance & Business ($\mathit{coef.}=0.072$, $\mathit{odds~ratio}=1.075$, $p<0.001$) are more likely to receive community notes compared to posts not associated with these topics. Additionally, posts containing media are substantially more likely to receive community notes compared to text-only posts ($\mathit{coef.}=0.205$, $\mathit{odds~ratio}=1.228$, $p<0.001$), corresponding to a 22.8% increase in the odds of receiving community notes. For claims vs. opinions, we find that posts that contain claims are associated with 14.5% higher odds of receiving a community note than posts without claims ($\mathit{coef.}=0.135$, $\mathit{odds~ratio}=1.145$, $p<0.001$), whereas posts expressing opinions are not significantly associated with the likelihood of receiving notes ($\mathit{coef.}=-0.034$, $\mathit{odds~ratio}=0.966$, $p=0.186$). Importantly, we find that posts with higher GPT-estimated misleadingness are significantly more likely to receive community notes ($\mathit{coef.}=0.257$, $\mathit{odds~ratio}=1.293$, $p<0.001$). Specifically, posts identified by GPT as purely misleading have 29.3% higher odds of receiving community notes than non-misleading posts.
@@
Post author characteristics. We further evaluate how post author characteristics shape the likelihood of receiving community notes. For verified types of post authors, we find that posts from blue ($\mathit{coef.}=-0.119$, $\mathit{odds~ratio}=0.888$, $p<0.001$) and business ($\mathit{coef.}=-0.205$, $\mathit{odds~ratio}=0.815$, $p<0.001$) accounts have 11.2% and 18.5% lower odds of receiving community notes than posts from unverified accounts, respectively. Additionally, posts from accounts with a higher number of followees are less likely to receive community notes, with one standard deviation increase in followers corresponding to 5.8% decrease in the odds of receiving notes ($\mathit{coef.}=-0.060$, $\mathit{odds~ratio}=0.942$, $p<0.001$). In contrast, misinformation exposure scores are positively associated with the likelihood of receiving community notes: posts from accounts with a misinformation exposure score of 1 have 88.3% higher odds than those from accounts with a score of 0 ($\mathit{coef.}=0.633$, $\mathit{odds~ratio}=1.883$, $p<0.001$). Moreover, posts from right-leaning accounts with a partisan score of 1 have 11.4% higher odds of being noted compared to those from left-leaning accounts with a partisan score of 1-1 ($\mathit{coef.}=0.056$, $\mathit{odds~ratio}=1.057$, $p<0.01$).
@@
In summary, the likelihood that a requested post receives a community note is shaped by both posts’ and their authors’ characteristics. Requested posts with negative sentiment or authored by blue/business accounts are less likely to attract notes, as are those related to politics or science. In contrast, posts related to entertainment or finance, those containing factual claims and media, and posts flagged by GPT as more misleading are significantly more likely to receive notes. At the author level, posts from accounts with a higher misinformation exposure or right-leaning partisanship (vs. left-leaning) are also more likely to be annotated. Notably, given the heightened concerns surrounding political misinformation and the fact that political posts constitute the largest part of our dataset, we additionally restrict our analysis to political posts and find that the observed patterns remain robust (see estimation results in Suppl. S4). Together, our findings highlight that Community Notes contributors who write notes and requestors who submit requests may have different selection patterns that are associated with both post features and author attributes.
@@ 4.2. Request Timing and Request-Fostered Notes
Request timing relative to note creation. As shown in Fig. 3(a), the median time from post creation to the submission of first request is 2.72 hours, while it takes significantly longer for a post to accumulate five requests (median of 13.40 hours; $p_{MWU}<0.001$). We find that the response time of community notes (median of 8.23 hours) is significantly shorter than the time it typically takes for posts to receive five requests (Fig. 3(b); $p_{MWU}<0.001$). Furthermore, Fig. 3(c) shows that top writers (median of 7.78 hours) generate notes almost an hour faster than other contributors (median of 8.76 hours; $p_{MWU}<0.001$). Taken together, these findings suggest that community notes tend to be written before the fifth request is submitted (i. e., request threshold)—implying that requests may not be the primary driver of note creation. Additionally, because community notes often appear more quickly than request alerts, and because top writers tend to response faster than other contributors, notes generated by top writers after the fifth request to the corresponding posts are highly likely to have been shaped by request alerts. Therefore, we consider top-writer notes generated after the request threshold as request-fostered notes. Consistent with the guidelines of the request feature, we find that only an estimated 12.1% of requested posts were likely influenced by the request activity, as their associated notes were written by top writers after the request threshold was reached (Fig. 3(d)). Therefore, the effect of the request feature in driving note creation is still limited.
@@
Request-fostered notes from top writers. We further examine whether the request function influences the fact-checking behavior of top writers. To this end, we employ the logistic regression model as specified in Eq. 1 and compare top writers’ selection patterns when exposed to request alerts versus when writing notes without such alerts. Here, the dependent variable is the probability that a post receives a request-fostered note. Under the null hypothesis that requests do not affect selection behavior of top writers, the coefficient estimates for all independent variables should be statistically indistinguishable from zero; statistically significant coefficient estimates, by contrast, indicate systematic differences in post selection following request exposure. Importantly, to ensure a fair comparison, we restrict our analysis to posts on which notes were generated only by topic writers before and after the request threshold. The estimation results are presented in Fig. 4.
We find that top-writer notes on posts related to Politics ($\mathit{coef.}=0.253$, $\mathit{odds~ratio}=1.287$, $p<0.001$) and Science & Technology ($\mathit{coef.}=0.091$, $\mathit{odds~ratio}=1.095$, $p<0.05$) have 28.7% and 9.5% higher odds of being request-fostered, respectively, compared to posts without the corresponding topics. This suggests that top writers select more political and scientific posts when receiving request alerts. In contrast, top-writer notes on posts related to Entertainment ($\mathit{coef.}=-0.095$, $\mathit{odds~ratio}=0.910$, $p<0.01$) and containing claims ($\mathit{coef.}=-0.138$, $\mathit{odds~ratio}=0.871$, $p<0.05$) have 9% and 12.9% lower odds of being request-fostered, respectively. This indicates that top writers choose less entertainment-related and claim-oriented content after receiving request alerts, compared to without receiving alerts. Together, these results suggest that requests can shift contributors’ attention toward political and scientific content, while diverting attention away from claim-oriented posts or those related to entertainment. In particular, the coefficient estimate for GPT-estimated misleadingness is positive and substantially significant ($\mathit{coef.}=0.489$, $\mathit{odds~ratio}=1.631$, $p<0.001$). Posts flagged by GPT as misleading (score =1=1) have 63.1% higher odds of receiving request-fostered top-writer notes than non-misleading posts (score =0=0). This suggests that top writers prioritize posts with higher potential misleadingness when guided by request alerts, compared to when no alerts are present.
In summary, out of all requested posts eligible for community evaluation, only 12.1% appear to have been generated by top writers through request alerts. Furthermore, requests typically arrive later than community notes themselves, as the time from post creation to request alerts is significantly longer than the time from post creation to note generation, limiting the request feature’s potential to foster note writing. Nevertheless, by comparing notes written by top writers before and after the request threshold, we find that requests shift contributors’ attention toward political content while prioritizing posts with higher potential misleadingness. [^4] To further reduce unobserved confounding at the contributor level, including heterogeneity among top writers, we additionally restrict the analysis to posts selected by the same set of top writers before and after the request threshold (see details in Suppl. S5). This conservative approach confirms that our main findings remain robust, particularly with respect to political content and posts with high potential misleadingness. Thus, although the overall effect of requests on note creation remains modest, the feature shows promise in channeling contributor attention toward high-stakes content at greater risk of misinformation.
@@ 4.3. Algorithmic Evaluation of Community Notes
To address RQ3, we downloaded and rerun the source code of the note selection algorithm released by $\mathbb{X}$. Fig. 5(a) presents the distribution of note intercepts and note factors from the note selection algorithm. The note intercepts reflect the perceived helpfulness of community notes across raters, with higher intercepts indicating higher helpfulness. The note factors capture the polarization in ratings, with values deviating from 0 indicating greater disagreement among raters regarding the helpfulness of community notes. Here, we categorize all community notes into three groups: writer-only (notes on posts that received fewer than five requests and were solely written by contributors), request-fostered (notes authored by top writers after the fifth request, likely influenced by requests), and request-related (all remaining notes on posts that received five or more requests but not classified as “request-fostered”). Notably, request-related notes may also be written by top writers; however, unlike request-fostered notes, they were produced before the threshold of request alerts. We then analyze the helpfulness and polarization (measured as the absolute value of note factors) of community notes across these categories.
@@
Figure 5. Overview of note evaluations and source domains in community notes. (a) The distribution of estimated note intercepts (helpfulness) and estimated note factors (polarization) from the note selection algorithm. (b) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and writer-only notes. (c) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-fostered notes and writer-only notes. (d) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and request-fostered notes. (e) The violin plots showing the distributions of domain bias in community notes across the three categories: writer-only, request-related, and request-fostered. (f) The violin plots showing the distributions of domain quality in community notes across the three categories: writer-only, request-related, and request-fostered.
@@
Overview of note evaluations and source domains in community notes. (a) The distribution of estimated note intercepts (helpfulness) and estimated note factors (polarization) from the note selection algorithm. (b) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and writer-only notes. (c) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-fostered notes and writer-only notes. (d) The distributions of estimated note intercepts (helpfulness) and note factors (polarization) from the note selection algorithm between request-related notes and request-fostered notes. (e) The violin plots showing the distributions of domain bias in community notes across the three categories: writer-only, request-related, and request-fostered. (f) The violin plots showing the distributions of domain quality in community notes across the three categories: writer-only, request-related, and request-fostered.
@@
Taken together, these findings indicate that, although requests in general are associated with less helpful and more polarized notes, request-fostered notes written by top writers tend to achieve higher quality—being both more helpful and less polarized than request-related notes. To further substantiate this pattern, we estimate a logistic regression predicting whether a note is request-fostered versus request-related using helpfulness and polarization as predictors. The results show a strong positive association for helpfulness ($\mathit{coef.}=2.528$, $\mathit{odds~ratio}=12.530$, $p<0.001$) and a significant negative association for polarization ($\mathit{coef.}=-0.231$, $\mathit{odds~ratio}=0.794$, $p<0.001$), confirming that request-fostered notes stand out within the request context by being both more helpful and less polarized than other notes. Subsequently, we investigate what factors could explain this quality advantage.
@@
Because Community Notes contributors are anonymized within the system, rater bias and polarization toward specific writers’ identities (e. g., partisanship) are reduced (Notes, 2025f). Consequently, observed differences in helpfulness or polarization are more likely to reflect the types of posts and authors contributors choose to fact-check. We therefore examine which notes, posts, and authors are associated with higher helpfulness or polarization.
Factors associated with note helpfulness. The estimation results from the linear regression model for note helpfulness (as specified in Eq. 2) are shown in Fig. 6. We find that the unbiasedness (i. e., neither left-leaning nor right-leaning) of external domains cited in community notes is positively associated with note helpfulness ($\mathit{coef.}=0.005$, $p<0.05$), suggesting that notes referencing politically neutral sources are perceived as more helpful compared to those citing left- or right-leaning domains. In contrast, the information quality of external domains is not significantly associated with helpfulness ($\mathit{coef.}=0.015$, $p=0.095$).
Regarding post content, notes on posts with higher negative sentiment are rated as less helpful ($\mathit{coef.}=-0.016$, $p<0.01$). Posts related to Politics ($\mathit{coef.}=-0.015$, $p<0.001$) and Science & Technology ($\mathit{coef.}=-0.012$, $p<0.001$) tend to receive less-helpful notes compared to posts without these topics. Additionally, posts with higher opinion scores are associated with lower note helpfulness ($\mathit{coef.}=-0.017$, $p<0.001$). Conversely, community notes on posts containing media ($\mathit{coef.}=0.009$, $p<0.001$) or with higher estimated misleadingness ($\mathit{coef.}=0.014$, $p<0.01$) are rated as more helpful. With respect to post authors, notes on posts from the blue ($\mathit{coef.}=-0.010$, $p<0.01$) and government ($\mathit{coef.}=-0.013$, $p<0.01$) accounts, as well as from authors with more followers ($\mathit{coef.}=-0.027$, $p<0.001$), are less helpful compared to those from unverified accounts with fewer followers. In contrast, higher misinformation exposure scores of post authors are positively associated with note helpfulness ($\mathit{coef.}=0.090$, $p<0.001$).
In summary, the helpfulness of community notes is positively associated with (i) citations to unbiased domains, (ii) posts containing media, (iii) post misleadingness, and (iv) misinformation exposure scores of post authors. Conversely, the helpfulness of community notes is negatively associated with (i) negative sentiment in posts, (ii) political and scientific topics, (iii) opinionated content, (iv) blue and government accounts, and (v) the number of followers.